OpenStudy (anonymous):

What is the greatest common factor of 40 and 24?

OpenStudy (zzr0ck3r):

what is the biggest number that divides them both?

OpenStudy (zzr0ck3r):

what are all the numbers that divide 24?

OpenStudy (anonymous):

12

OpenStudy (zzr0ck3r):

does 12 divide 40?

OpenStudy (zzr0ck3r):

here I will help you to show you the general way to solve these problems. pick the smallest number 24 list its divisors 1,2,3,4,6,8,12,24 (bonus: The fact that it has this many divisors is part of the reason we use 24 for the amount of hours in a day.) OK start at the top Now 24 does not divide 40 because 24*2 = 48 Does 12 divide 40? 12*2 = 24 12*3 = 36 12*4 = 48 It does not, so we move to the next. Does 8 divide 40?

OpenStudy (legends):

Psst the answer is 4.